A pottery wheel with a mass of 10kg and a radius of .15m is being used to shape some clay. If the wheel is initially spinning with an angular speed of 12.87 rad/s and it is subject to a constant angular acceleration for 7.66 until it stops. What is the magnitude of angular acceleration of the wheel?
Answer #1

Angular acceleration,

a = w/t = 12.87/ 7.66

a = 1.68 rad/sec^2

Comment in case any doubt please rate my answer ....